Use variation of parameters to find the general solution of the differential equation: đư” =T.

please show all work
Answer #1

Given differential eq? is 44 = x² or y = x²/4 Its auxiliary eqn. will be m=0 m=0,0 Thus, Y c = (6 + x) 8.x = (G+CqX) 1 = t